# Supermassive black hole growth from stellar binary encounters

### v2

@@ 2.2 A channel for black hole growth
Every interaction between a binary star and a supermassive black hole that leads to binary break-up places a star on a bound orbit that will eventually lead to tidal disruption as a result of orbital diffusion (e.g., Bromley et al., 2012). While tidal disruption outside the horizon may not result in accretion of all the freed stellar gas, at least initially (Evans & Kochanek, 1989), we assume that for stars already on bound orbits, all the disrupted gas ultimately is accreted by the SMBH. In this scenario, the binary break-up rate in Equation (6) directly provides the black hole growth rate,
@@
  • The black hole mass $M_{\bullet}$ is a function of time $t$, with a initial mass $M_{\bullet}(0)$ at $t=0$, the present epoch.
  • The central stellar density $\rho_{0}$ is also a function of time through its dependence on $M_{\bullet}(t)$, wherein $\rho_{0}\sim M_{\bullet}/(M_{b}^{2}+M_{\bullet}^{2})$ (cf. Eq. (4)), scaled so that $\rho_{0}(0)$ is the present-day value.
@@
where $A$ is a constant (cf. Eq. (7)). Its solution, from separation of variables, is
@@
where $X=At+M_{\bullet}(0)-M_{b}^{2}/M_{\bullet}(0)$ and $M_{\bullet}(0)$ is the black hole mass at $t=0$, the present epoch.
Here, we also use an explicit, step-wise numerical integrator (the Euler method), written in Python, to solve for black hole growth. We start with the initial SMBH mass $M_{\bullet}(0)$, evaluate all parameters ($\rho_{0}$, $f_{\text{grav}}$, $v_{\rm disp}$, etc.) at that time, and calculate $km_{\rm avg}$. The mass at a time $t=\Delta t$ later is $M_{\bullet}(0)+km_{\rm avg}\Delta t$. This process is repeated to step through to some desired ending time. In practice we use $\Delta t=0.1$–1 Myr. This numerical approach allows us to explore more complicated parameterizations and growth scenarios. Figure 1 provides an illustration.
Figure 1: The evolution of black hole mass from the Hills mechanism and stellar accretion. Each curve shows a unique starting mass, which we evolve usng the growth equation (Eq. (10)). We adopt the analytical relationship between black hole mass and central stellar density in Equation (4), while all other parameters are assigned their fiducial values as in Equation (7). Because the central density is near a maximum for $_{\bullet}\approx 4\times 10^{7} M\rm M_{\odot}$, the most rapid growth occurs for black holes around this size.
@@ 3 SMBH growth in nearby galactic hosts
We apply our binary break-up model to observed galactic sources to evaluate the potential growth of supermassive black holes from binary break-up and stellar accretion under varying conditions. While many of the parameters of our model are unknown, estimates are available for the two critical ones, SMBH mass $M_{\bullet}$ and the central stellar density $\rho_{0}$. Here, we consider black hole growth using a curated dataset from Hannah et al. (2024), consisting of 91 galaxies of a range of types, all within 50 Mpc. With the Python Pandas package, we selected data by filtering to include only black holes with mass derivation methods corresponding to source codes 1, 2, and 3, ensuring methodological consistency. We thus identified 30 well-measured galaxies to use in this study. The remaining 61 sources have black hole masses that are not well constrained. For each selected galaxy, we extracted the galaxy name, the black hole mass (in logarithmic solar mass units) and central stellar density (in logarithmic solar masses per cubic parsec), defined as an average inside a distance of 5 pc from the dynamical center of the galaxy.
Figure 2 illustrates the dataset, showing all of the sources in Hannah et al. (2024), including measurements not considered in this study. The plot also shows our analytical expression that broadly tracks the central density as a function of black hole mass (Eq. (4)).
Figure 2: Central density versus SMBH mass. The larger, violet points correspond to sources with more reliable black hole mass estimates, while the smaller, turquoise points represent galaxies with poorly constrained black hole masses. The gray line is a simple analytical model for the mass-density relation (Eq. (4)). The low-mass end of this model is speculative.
To assess the potential for growth, we start with the estimated SMBH mass of each galaxy and the reported central stellar density. We then integrate the mass forward in time as in Figure 1, with an important difference: The central stellar density $\rho_{0}$ is given the same form of dependence on black hole mass as in Equation (4), but the normalization is set to match the density in Hannah et al. (2024). Our results are presented in Figure 3, which shows predicted future growth of the 30 galaxies that we selected from the Hannah et al. (2024) dataset.
@@
Another view of black hole growth comes from integrating the growth equation (Eq. (10)) backward in time to assess the fraction of the observed black hole mass may have arisen from binary break-ups. In this way, we predict the fraction of a black hole’s mass that could potentially have come from the Hills mechanism and stellar accretion. Figure 4 provides that view.
@@
Figure 4: Fraction of the black hole mass predicted to have come from stellar accretion through the Hills mechanism. The diamond symbols are the predictions for the 30 galaxies with high-quality SMBH masses in the (Hannah et al., 2024), run backwards in time for 1 Gyr with our numerical integrator. The short vertical lines near the symbols are the predicted fraction under the assumption that the central density as been fixed in time. For black holes smaller than $_{b}\sim 4\times 10^{7} M\rm M_{\odot}$, the constant-density models overpredict this fraction, since the stellar density tends to decrease with decreasing black hole mass. For black holes with mass larger than $M_{b}$, the constant-density models underpredict the fraction of mass from stellar accretion.
The predicted mass evolution in Figures 3 and 4 illustrates that conditions for black hole growth vary widely among galaxies. Of the 30 sources we consider here, 20% (6) are predicted to have at least doubled the black hole mass in the past gigayear, while in 8 galaxies, stellar accretion from binary break-ups accounts for less than 10% of the black hole mass.
The evolution calculations highlight two galaxies that exhibit exceptional SMBH growth rates over gigayear timescales: NGC 4342 and NGC 4026. The central SMBH in NGC 4342, a lenticular (S0) galaxy, in our calculations has formally acquired over 99% of its mass from the binary-capture mechanism over the last gigayear, and is predicted to grow by another order of magnitude over the next 5 Gyr. However, our predictions likely overestimate the potential of the stellar accretion growth channel. First, the measured velocity dispersion is significantly higher than the 100 km/s value in our model, which would mean we overestimate the binary break-up rate because our gravitational focusing factor is too high. Cretton & van den Bosch (1999) report $v_{\rm disp}\approx 400$ km/s from the Hubble Space Telescope. Second, the binary fraction in this early-type galaxy, presumably composed of lower-mass, older stars, is likely lower than our fiducial value of 0.5. Adopting conservative values of $v_{\rm disp}=400$ km/s and $f_{\rm bin}=0.1$ (see Basri & Reiners, 2006, for example), our model still predicts that over 20% of the SMBH mass in NCG 4342 came from this growth channel in the last gigayear, and more than 95% over the past 5 Gyr.
Similarly, the SMBH in NGC 4026, also an early-type S0 galaxy, evolves from an initial mass of $1.82\times 10^{8}$ $\rm M_{\odot}$ to over $10^{9}\,M_{\odot}$ within a couple of gigayears in our model. The fraction of mass from binary capture is formally over 95% (Fig. 4). Like the case of NGC 4342, this galaxy has a measured velocity dispersion that is greater than our fiducial value. Adopting a velocity disperson of 250 km/s (Gültekin et al., 2009), and the lower value of $f_{\rm bin}=0.1$, we find that binary captures have contributed to about 20% of the present-day SMBH mass in the past gigayear, and over 80% of its mass in the past 5 Gyr.
@@
We conclude that for the population of galaxies considered here, the binary-capture channel can play a significant, and possibly even dominant role in black hole growth.
@@ 4 Closer to home: the case of LMC*
Han et al. (2025) recently proposed the existence of a $6\times 10^{5}$ $\rm M_{\odot}$ black hole within the Large Magellanic Cloud (LMC). Their inference is based on the locations and velocities of the observed HVSs from the HVS Survey (Brown et al., 2012). This analysis incorporated Gaia Data Release 3 astrometry (Gaia Collaboration et al., 2023) to reveal that half of the HVS Survey sources originated in the LMC. With forward modeling using the ejection speed parameterization from Bromley et al. (2006) and an HVS production rate of $k=2$ Myr-1, Han et al. (2025) reproduced the observed phase-space distribution of these HVSs, including the velocity boost from LMC motion in the Galactic frame and the observed clustering on the sky within the context of the survey selection. The success of this modeling requires the presence of LMC.
@@
The results in Han et al. (2025) have implications for the growth of the dwarf galaxy’s massive black hole. The observed cluster of nine main-sequence, B-type HVSs linked to an LMC origin, found in the constellation Leo, turn out to be the tip of an iceberg representing binary star encounters with LMC. First, these stars are at the edge of the HVS Survey window, which do not include a large swath of sky around the LMC where the majority of HVSs are expected to reside. From the reported rate of HVS production required to match observations, these nine stars are associated with $\sim$800 other ejected stars with similar masses that lie beyond the surveyed sky. Second, B-stars are comparatively rare, and are the tip of another iceberg from a stellar population perspective. With a Salpeter initial mass function (IMF), stars with mass $m\geq 2$ $\rm M_{\odot}$ — the range considered by Han et al. (2025) — represent less than 2% of all stars with masses greater than 0.1 $\rm M_{\odot}$. The nine HVSs from LMC thus suggest a binary disruption rate greater than $10^{-4}$ yr-1. With an average mass around $0.3$ $\rm M_{\odot}$ captured onto tight orbit around LMC per disruption, the SMBH can acquire a significant fraction of its mass over a few gigayears.
@@
There are significant uncertainties in this growth rate. A major source is the initial mass function, since the HVS Survey only sampled the most massive stars. A top-heavy IMF would reduce the rate we infer here — the observed stars then are the tip of a less significant iceberg. Other sources of uncertainty include the details that affect the Hills mechanism and its outcome, including the distribution of binary orbits, the periastron distribution (empty or full loss cone Lightman & Shapiro, 1977; Stone & Metzger, 2016), and the mass distribution around the SMBH that impacts the speed (and interpretation) of the HVS candidates (e.g., Bromley et al., 2006; Sari et al., 2010; Rossi et al., 2014; Han et al., 2025).
A check on our overall approach (§2.2) is a comparison between our predicted binary break-up rate and the Han et al. (2025) LMC HVS study. With a velocity dispersion of 35 km/s, characteristic of the central region of the LMC (Vasiliev, 2018), and (in absence of a compelling value from the literature) a central density of $\rho_{0}\approx 750$ $\rm M_{\odot}$/pc3 from Equation (4), we find $k=2.0\times 10^{-4}$, similar to the estimate based on the HVS number count reported in Han et al. (2025). Our model predicts that about 10% of LMC’s mass came from stellar accretion over the last gigayear, and 30% in the past 5 Gyr. Although binary accretion on its own cannot account for the emergence of a $10^{6}$ /msun/ black hole, the mechanism can help build a SMBH from a smaller seed object. A recent detailed study of black hole growth in globular clusters came to a similar conclusion (González Prieto et al., 2025).

### v3

@@ 2.2 A channel for black hole growth
Every interaction between a binary star and a supermassive black hole that leads to binary break-up places a star on a bound orbit that will eventually be captured by the SMBH as a result of orbital diffusion (e.g., Bromley et al., 2012). Although tidal disruption outside the horizon may not result in accretion of all the freed stellar gas, at least initially (Evans & Kochanek, 1989), we assume that for stars already on bound orbits, all the disrupted gas ultimately is accreted by the SMBH. In this scenario, the binary break-up rate in Equation (2.1) directly provides the black hole growth rate,
@@
  • The black hole mass $M_{\bullet}$ is a function of time $t$, with an initial mass $M_{\bullet}(0)$ at $t=0$, the present epoch.
  • The central stellar density $\rho_{0}$ is also a function of time through its dependence on $M_{\bullet}(t)$ in Equation (5), scaled so that $\rho_{0}(0)$ is the present-day value.
@@
where $A=km_{\rm avg}$ is a constant (cf. Eq. (LABEL:eq:kfid)). Its solution, from separation of variables, is
@@
where $X=At+M_{\bullet}(0)-M_{b}^{2}/M_{\bullet}(0)$ and $M_{\bullet}(0)$ is the black hole mass at $t=0$, the present epoch. This approximation gives a good representation of growth in most cases of interest here, wherein the SMBH mass increases by some fractional amount over the course of a gigayear.
We also use a step-wise numerical integrator (the Euler method), written in Python, to solve for black hole growth including the variation of all parameters with SMBH mass (Eq. 10). We start with the initial SMBH mass $M_{\bullet}(0)$ and calculate $km_{\rm avg}$. The mass at a time $t=\Delta t$ later is $M_{\bullet}(0)+km_{\rm avg}\Delta t$. This process is repeated until we reach some desired end time using $\Delta t=0.1$–1 Myr. This numerical approach allows us to explore more complicated parameterizations and growth scenarios. Figure 1 provides an illustration.
Figure 1: The evolution of black hole mass from the Hills mechanism and stellar accretion. Each curve shows a unique starting mass, which we numerically evolve using the growth equation (Eq. (12)). We adopt the analytical relationship between black hole mass and central stellar density in Equation (5), and include the mass-dependence of other parameters based on the fiducial values in Equation (LABEL:eq:kfid). The analytical solutions (Eq. (13), dashed curves) are shown for comparison. Because the central density is near a maximum for $_{\bullet}\approx 4\times 10^{7} M\rm M_{\odot}$, the most rapid growth occurs for black holes around this size.
2.3 A modified kinematic model
The rapid growth of a SMBH with masses greater than $_{b}\approx 10^{8} M\rm M_{\odot}$ predicted by the simple kinematical model in Equation (12) may be challenging to realize. Growth may be limited by “loss-cone” effects, wherein binary break-ups and the accretion of stars generally proceed at a much slower rate than predicted by simple kinematics by virtue of the depletion of orbits that pass close to the SMBH (Frank & Rees, 1976; Lightman & Shapiro, 1977; Cohn & Kulsrud, 1978; Magorrian & Tremaine, 1999; Perets et al., 2007; Merritt, 2013, see also Hannah et al. 2025 and references therein). Furthermore, even if phase-space mixing rapidly replenished orbits in the loss cone, the mass reservoir in stars surrounding the black hole may insufficient to affect its overall mass.
For a better assessment of growth rates in light of loss-cone effects and limited mass reservoirs surrounding SMBHs, we introduce two changes. First, we modify the stellar mass density in the rate estimation to the value not at a fixed distance (5 pc) from a SMBH, but at the outer limit of the black hole’s sphere of influence to represent unbound stars (Eq. (2)). Then, to enforce the depletion of orbits in loss cones around the most massive SMBHs, we introduce a factor of $1/(1+M_{\bullet}/M_{b})$. The result is an ansatz, a modified kinematical model,
$\frac{dM_{\bullet}}{dt}=AM_{b}r_{v}^{-\gamma}\frac{M_{\bullet}^{2+\gamma}}{(M_{\bullet}^{2}+M_{b}^{2})^{2}},$ (14)
where $A$ is the same constant as in Equation 12 and $\gamma$ is the power-law index for the stellar density profile (Eq. (6)). We validate this new parameterization by comparing with detailed TDE rates in §4. We next explore black hole growth in observed galaxies with both the kinematical model and this modified version.
@@ 3 SMBH growth in nearby galactic hosts
We apply our binary break-up models to observed galactic sources to evaluate the potential growth of supermassive black holes from binary break-up and stellar accretion under varying conditions. While many of the model parameters are unknown, estimates are available for the two critical ones, then SMBH mass $M_{\bullet}$ and the central stellar density $\rho_{0}$. Here, we consider black hole growth using the curated dataset from Hannah et al. (2024), consisting of 91 galaxies of a range of types, all within 50 Mpc. With the Python Pandas package, we highlight data by filtering to include black holes with mass derivation methods corresponding to source codes 1, 2, and 3, ensuring methodological consistency. We thus identify 30 well-measured galaxies to showcase in this study. The remaining 61 sources have black hole masses that are not as well constrained. For each selected galaxy, we extracted the galaxy name, the black hole mass (in logarithmic solar mass units) and central stellar density (in logarithmic solar masses per cubic parsec), defined as the value at a distance of 5 pc from the dynamical center of the galaxy.
Figure 2 illustrates the dataset, showing the representative central stellar mass density (evaluated at 5 pc from the galactic center) versus SMBH mass for all of the sources in Hannah et al. (2024, see also ). The plot also shows our analytical expression that broadly tracks the central density as a function of black hole mass (Eq. (5)).
Figure 2: Central density versus SMBH mass. The larger, violet points correspond to sources with more reliable black hole mass estimates, while the smaller, turquoise points represent galaxies with poorly constrained black hole masses. The gray line is a simple analytical model for the mass-density relation (Eq. (5)). The low-mass end of this model is speculative. In all cases, the stellar densities are evaluated at 5 pc from the galactic centers.
To assess the potential for growth, we start with the estimated SMBH mass of each galaxy and the reported central stellar mass density. We then integrate the mass forward in time as in Figure 1 using our kinematical model (Eq. (12)). The central stellar density $\rho_{0}$ is given the same form of dependence on black hole mass as in Equation (5), but the normalization is set to match the density in Hannah et al. (2025). Our results are presented in Figure 3, which shows predicted future growth of the 30 galaxies from the Hannah et al. (2024) dataset with the most reliable SMBH masses.
@@
Another view of black hole growth comes from integrating the growth equation from the simple kinematical ($nv\sigma$”) model (Eq. (12)) backward in time. In this way, we predict the fraction of a black hole’s mass that could potentially have come from the Hills mechanism and stellar accretion. Figure 4 provides that view. It also shows the predictions from the modified model with stellar density set to its value at the radius of influence of the black hole and a term to suppress growth of black holes with $M_{\bullet}\gtrsim M_{b}$ (Eq. (14)).
@@
Figure 4: Fraction of the black hole mass predicted to have come from stellar accretion through the Hills mechanism. The diamond symbols are the predictions for the 30 galaxies with high-quality SMBH masses in the Hannah et al. (2025) data set, run backwards in time for 1 Gyr with our numerical integrator.
The predicted mass evolution in Figures 3 and 4 illustrates that conditions for black hole growth vary widely among galaxies and our model assumptions. Of the 91 sources in the Hannah et al. (2024) data set, the SMBHs in 39 galaxies are predicted to have grown in mass by more than 10% in the last gigayear, according to the simple kinematical model. All of these galaxies have SMBHs with masses near or above $10^{7}$ $\rm M_{\odot}$. In the modified model, 37 galaxies have SMBHs that grew by that same amount in a gigayear, and most of these galaxies have lower-mass SMBHs. A total of 11 galaxies have SMBHs that have grown by 10% or more in both models. Table 1 lists these galaxies.
Table 1: Promising candidates for black hole growth by binary break-up
$\log_{10}(M_{\bullet})$ $\log_{10}(\rho_{0})$ growth [%] qual.
name $\rm M_{\odot}$] $\rm M_{\odot}$/pc3] $nv\sigma$ mod. flag
NGC 2778 7.150 4.290 33.28 35.76 1
NGC 3384 7.040 3.750 12.16 14.78 1
NGC 3522 7.100 3.730 11.82 12.36 0
NGC 3945 6.940 4.110 23.66 41.50 1
NGC 4387 7.160 3.750 12.48 11.24 0
NGC 4458 7.120 3.790 13.41 12.70 0
NGC 4483 7.160 3.870 15.85 14.60 0
NGC 4612 6.770 3.700 10.44 28.54 0
NGC 4638 7.510 4.370 42.50 21.11 0
NGC 5236 6.910 3.950 17.57 33.67 0
NGC 7457 6.950 3.790 12.95 22.00 1
[…]
While a few lower-mass SMBHs experience rapid growth (e.g., SMBHs in NGC 4605 and NGC 5238, with growth of over 90% in the past gigayear) in the modified kinematic model, and a couple of high-mass SMBHs (in NGC 4660 and NGC 4342) similarly stand out in the simple model, only the 11 galaxies in Table 1 have significant SMBH growth in both models. Of these, two — NGC 2778 and NGC 3945 — are standouts because they have reliable SMBH mass measurements ($_{\bullet}\sim 10^{7} M\rm M_{\odot}$) and SMBH growth rates above 20% per gigayear.
NGC 2778 is an elliptical galaxy with a central velocity dispersion of $v_{\rm disp}\approx 150$ km/s (Jin et al., 2025) and a mass $_{\bullet}=1.4\times 10^{7} M\rm M_{\odot}$. The growth rate in this case may be overstated because the stellar velocity dispersion is higher than our adopted value by a factor of about 40%. Given that the binary fraction we have adopted is modest ($f_{\rm bin}=0.1$), we still expect that this galaxy may be a promising candidate for the binary break-up growth channel.
The galaxy NGC 3945 has a barred lenticular morphology with prominent rings and a central velocity dispersion of about 170 km/s (Gültekin et al., 2009). While this dispersion is also high compared with our prescription given $_{\bullet}\approx 10^{7} M\rm M_{\odot}$, it still leads to a predicted growth rate of 13–24% per gigayear, depending on the kinematic model. Furthermore, the non-axisymmetric distribution of stars in this barred galaxy offers promise of phase-space mixing and perhaps loss-cone replenishment beyond the SMBH’s sphere of influence (e.g., Sellwood & Wilkinson, 1993).
@@
We conclude that for many of the galaxies considered here, the binary-capture channel may play a significant, and, in some cases, even a dominant role in black hole growth.
4 Comparison with predictions of tidal disruption events
Our estimates of SMBH growth from binary break-up are based on encounter rates from simple $nv\sigma$” kinematics, a choice that highlights what might be possible, not necessarily what is plausible. Hannah et al. (2025) performed more detailed calculations for stellar encounters that lead to TDEs for the same set of galaxies considered here. They used a sophisticated phase-space evolution code, REPTiDE (Hannah, 2025), with inputs that include the stellar mass density profile (density at 5 pc and the power-law index $\gamma$) and black hole masses. Their published TDE rates offer an opportunity for us to connect the impact of assumptions made here for binary break-up.
We compare our approach with REPTiDE results by adapting our simple and modified kinetic models to stellar disruption instead of binary break-up. In the event rate estimate (Eq. (2.1)), we remove dependence on $f_{\rm bin}$ and replace $r_{\rm close}$ with a representative stellar tidal radius
$r_{\rm tid}=r_{\ast}\left(\frac{m_{\ast}}{M_{\bullet}}\right)^{1/3}$ (15)
where $m_{\ast}$ is the star’s mass (here, $m_{\ast}=m_{\rm avg}=0.3$ $\rm M_{\odot}$) and $r_{\ast}\propto m_{\ast}^{0.8}$ is its radius.
Figure 5 is the result. The trend in the REPTiDE predictions is broadly decreasing TDE rates with SMBH mass. Hannah et al. (2025) note that the stellar phase-space distribution for galaxies with smaller SMBH masses have “full loss cones”; there is little depletion of radial orbits that can take stars close to the SMBH. For higher-mass black holes, the loss-cone becomes depleted. In our simple kinematic model, we assume a full loss cone in all cases. Thus, our predictions of TDE rates for the highest-mass black holes are relatively high. For lower-mass SMBHs, our model predicts a comparatively low rate. The reason is that the stellar density in our simple model is assessed at 5 pc, well beyond the SMBH radius of influence, and likely an underestimate of the density of unbound stars that interact with the SMBH.
Figure 5: Tidal disruption event rates from REPTiDE (Hannah et al., 2025) and the two kinematic models presented here, adapted to encounters between single stars and a SMBH. The black symbols are the predictions from REPTiDE, while the dark orange symbols are from our simple kinematic model. Our model overestimates the TDE rate at large SMBH mass, and underestimates it at low mass. The modified kinematic model predictions (light blue symbols) better track the REPTiDE estimates.
In contrast, our modified kinematical model, which uses an estimate of the stellar mass density just outside the sphere of influence, broadly matches the REPTiDE results for SMBH masses up to $_{\bullet}\sim 10^{9} M\rm M_{\odot}$. Indeed, our parameterization in Equation (14) was guided by this comparison.
The success of our modified kinematic model in reproducing TDE rates from loss-cone theory calculations highlights a liability. Binary break-up rates can be significantly higher than loss-cone predictions because the interaction cross section depends on the orbital separation, which is much larger than the tidal radius of main-sequence stars. Yet even loss-cone theory for the capture of single stars may over-predict observed TDE rates. For example, Yao et al. (2023) analyzed TDEs observed in optical light and concluded that TDE event rates per galaxy were well below $10^{-4}$ per year. Low-number statistics, waveband selection, and flux limits (Hannah et al., 2025, §4 therein), as well as accounting for partial TDEs (e.g., Bortolas et al., 2023), may all play a role in resolving this tension between theory and observation. However, an increase in the predicted rate of TDEs from captured binary partners on tight orbits may be challenging.
We offer the following considerations that might support the role of binary break-up in SMBH growth. Black holes with mass greater than the Hills mass,
$M_{{}_{\rm Hills}}=\left(\frac{c^{2}r_{\ast}}{2Gm_{\ast}^{1/3}}\right)^{3/2},$ (16)
will accrete solar-type stars without an observable TDE, since tidal disruption occurs inside the horizon (Hills, 1975). For stars of mass $m_{\rm avg}$, $M_{{}_{\rm Hills}}$ turns out to be close to $M_{b}$. If the predictions of the simple kinematic model (binary star-SMBH encounter rates increase with black hole mass) have relevance to these larger black holes, then binary break-up is an avenue for depositing stars deep inside a SMBH’s sphere of influence (Eqs. (2) and (2)) irrespective of observed constraints on TDE rates. Evidence for this potential growth channel would necessarily be indirect, possibly from a gravity wave signature (e.g., Toscani et al., 2025).
Lower-mass black holes with $_{\bullet}\lesssim 10^{6} M\rm M_{\odot}$ (including intermediate-mass black holes) can experience significant growth even with the low TDE rates inferred from observations (Milosavljević et al., 2006; Stone & Metzger, 2016). Binary break-up may be a promising avenue to bring stars close enough a lower-mass central black hole to be fully disrupted and accreted. We explore a compelling example, next, wherein binary break-up is directly implicated.
@@ 5 Closer to home: the case of LMC*
Han et al. (2025) recently proposed the existence of a $6\times 10^{5}$ $\rm M_{\odot}$ black hole within the Large Magellanic Cloud (LMC). Their inference is based on the locations and velocities of the observed HVSs from the HVS Survey (Brown et al., 2012). This analysis incorporated Gaia Data Release 3 astrometry (Gaia Collaboration et al., 2023) to reveal that half of the HVS Survey sources originated in the LMC. With forward modeling using the parameterization of the ejection speed of Bromley et al. (2006) and an HVSs production rate of $k=2$ Myr-1, Han et al. (2025) reproduced the observed phase-space distribution of these HVS, including the velocity increase from LMC motion in the Galactic frame and the observed clustering in the sky within the context of the survey selection. The success of this modeling requires the presence of LMC.
@@
The results in Han et al. (2025) have implications for the growth of the dwarf galaxy’s massive black hole. The observed cluster of nine main-sequence, B-type HVSs linked to an LMC origin, found in the constellation Leo, turn out to be the tip of an iceberg representing binary star encounters with LMC. First, these stars are at the edge of the HVS Survey window, which do not include a large swath of sky around the LMC where the majority of HVSs are expected to reside. From the reported rate of HVS production required to match observations, these nine stars are associated with $\sim$800 other ejected stars with similar masses that lie beyond the surveyed sky. Second, B-stars are comparatively rare, and are the tip of another iceberg from a stellar population perspective. With a Salpeter initial mass function (IMF), stars with mass $m\geq 2$ $\rm M_{\odot}$ — the range considered by Han et al. (2025) — represent less than 2% of all stars with masses greater than 0.1 $\rm M_{\odot}$. The nine HVSs from LMC therefore suggest a binary break-up rate greater than $10^{-4}$ yr-1. With an average mass around $0.3$ $\rm M_{\odot}$ captured on tight orbit around LMC per break-up, the SMBH can acquire a significant fraction of its mass over a few gigayears.
@@
There are important uncertainties in this growth rate. A major source is the initial mass function, since the HVS Survey only sampled the most massive stars. A top-heavy IMF would reduce the rate we infer here — the observed stars then are the tip of a less significant iceberg. Other sources of uncertainty include details that affect the Hills mechanism and its outcome, including the distribution of binary orbits, the periastron distribution (empty or full loss cone Lightman & Shapiro, 1977; Stone & Metzger, 2016), and the mass distribution around the SMBH that impacts the speed (and interpretation) of the HVS candidates (e.g., Gould & Quillen, 2003; Bromley et al., 2006; Sari et al., 2010; Rossi et al., 2014; Han et al., 2025).
A check on our overall approach (§2.2) is a comparison between our predicted binary break-up rate and the Han et al. (2025) LMC HVS study. We adopt a velocity dispersion of 35 km/s, characteristic of the central region of the LMC (Vasiliev, 2018), a central density at 5 pc of $\rho_{0}\approx 750$ $\rm M_{\odot}$/pc3 from Equation (5), and our fiducial binary fraction of $f_{\rm bin}=0.1$ (likely an underestimate, cf. Dalton & Sarazin, 1995). Absent a measurement of a stellar nucleus, we use $\gamma=-7/4$ to extrapolate $\rho$ inward from 5 pc (Eq. (6)) in our modified kinematic model. Our predicted binary break-up rates in the simple and modified models are $k=4\times 10^{-5}$ yr-1 (simple) to $k=2\times 10^{-4}$ yr-1 (modified), in the range of an estimate based on the HVS number count reported in Han et al. (2025). At these rates, between 10% and 40% of LMC’s mass came from stellar accretion over the past five gigayears. A modest increase in the assumed binary fraction would correspond to an even greater role of binary break-up for the black hole’s growth.
Although binary accretion on its own cannot account for the emergence of a $10^{6}$ $\rm M_{\odot}$ black hole, the mechanism can help build a SMBH from a smaller seed object. A similar result was found in a recent detailed study of black hole growth in globular clusters (González Prieto et al., 2025).